# Energy-Efficient Approximate Full Adders Applying Memristive Serial IMPLY Logic For Image Processingthanks: Preprint Submitted to arXiv

### v1

@@ 3 Proposed IMPLY-based approximate full adders
In this section, the proposed circuits and their design method are described. In subsection 3.1, three approximate full adders are introduced to reduce the hardware complexity along with an acceptable decrease in computation accuracy. In subsection 3.2, the primary purpose of proposing an approximate full adder is to avoid inexact carry propagation from LSBs to MSBs in an n-bit approximate adder structure.
@@ 3.1 Inexact Carry, Inexact Sum IMPLY-based approximate full adders (ICIS)
Three ICIS full adders (ICIS1-3) are designed in three steps and based on redesigning approximate logic from exact logic by changing the exact full adder’s truth table. It is necessary to determine the conditions of acceptability of approximate full adders’ accuracy before explaining the design steps of the proposed circuits. The maximum value of $ER_{Sum}$ and $ER_{C_{out}}$ should be $\frac{3}{8}$ and $\frac{1}{8}$ to have acceptable output results based on the error analysis metrics of SOA [3, 4, 14, 15, 16, 18, 29] in error-tolerant applications, e.g., image processing. According to these ERs, the maximum acceptable value of ED for the design of the proposed approximate full adders is also considered 3. Determining these conditions for designing ICIS1-3 is done because the reduction of computation’s accuracy should be limited. In addition to the error evaluation criteria, the computational steps should be significantly reduced compared to the exact full adders. The number of computational steps affects energy consumption directly.
The design steps of ICIS1-3, according to the mentioned design constraints, are:
@@
STEP 1: First, the truth table of the exact full adder is assessed (see Table 4). Then, according to the design constraints mentioned above, the $C_{out}$ of the exact full adder is inverted only in the first state ($A_{in}B_{in}C_{in}$=“000”), and $ER_{C_{out}}$ is $\frac{1}{8}$. In IMPLY logic, it is possible to invert an output in only two computational steps. In the first step, a memristor resets; in the second cycle, IMPLY function performs between the output memristor and the one reset in the last cycle. The $Sum$ output of these approximate full adders is assumed to be $\overline{C_{out}}$ to reduce the hardware complexity and the number of computational steps. This process is repeated for states 2 ($A_{in}B_{in}C_{in}$=“001”) to state 8 ($A_{in}B_{in}C_{in}$=“111”) of the exact full adder. Eight approximate full adders’ truth tables are designed by applying this method. We call these eight Approximate Full Adders AFA1-8. These eight approximate full adders’ truth tables can be seen next to the exact full adder’s truth table in Table 4. The truth tables of approximate full adders display exact outputs with ✓ and inexact ones with a ✕. ED of each AFA is specified in Table 4. $ER_{C_{out}}$ of these eight approximate full adders is $\frac{1}{8}$. The $ER_{Sum}$ of AFA2-7 is $\frac{3}{8}$, and the $ER_{Sum}$ of AFA1 and AFA8 equals $\frac{1}{8}$. It should be noted that the ED of all these approximate full adders equals 3. Among the AFA1-8 (see Table 4), the truth tables of AFA4, AFA6, and AFA7 are similar to those of SIAFA3, SIAFA1, and SIAFA4 introduced and implemented in our previous research [3]. So, AFA1-3, 5, and 8 are considered implementation candidates according to the error analysis metrics, and the truth table of these approximate full adders is not the same as the truth tables of SIAFA1, SIAFA3, and SIAFA4 [3].
@@
$Sum$ $C_{out}$ $Sum$ $C_{out}$ $Sum$ $C_{out}$ $Sum$ $C_{out}$ $Sum$ $C_{out}$ $Sum$ $C_{out}$ $Sum$ $C_{out}$ $Sum$ $C_{out}$ $Sum$ $C_{out}$
@@
$Sum$ $C_{out}$ $Sum$ $C_{out}$ $Sum$ $C_{out}$ $Sum$ $C_{out}$ $Sum$ $C_{out}$ $Sum$ $C_{out}$ $Sum$ $C_{out}$ $Sum$ $C_{out}$
@@
This time the truth table of eight approximate full adders is generated by inverting one bit of the exact full adder’s $Sum$ output from state $A_{in}B_{in}C_{in}$=“000” to state $A_{in}B_{in}C_{in}$=“111” similar to what was done for exact $C_{out}$. This time, $C_{out}$ equals $\overline{Sum}$, and $C_{out}$ can be calculated with only two computational steps. The truth tables of Approximate Full Adders 9-16, AFA9-16, are tabulated in Table 5. According to this design method, the $ER_{Sum}$ of these approximate full adders equals $\frac{1}{8}$. As in Table 4, the ED of AFA9-16 is written in the last row of Table 5. AFA9 and AFA16 have acceptable error analysis metrics (See Table 5) based on the conditions specified about the ER and ED of approximate full adders. According to the truth tables of these two circuits, it can be concluded that AFA9 and AFA16 are logically equivalent to AFA1 and AFA8.
STEP 2: After determining the truth table of the acceptable circuits from STEP 1, the boolean logic function of the outputs of each circuit must be determined. Based on these functions, an algorithm can be provided for implementing these circuits with IMPLY logic by applying the IMPLY and FALSE functions to memristors. Karnaugh map is one of the methods that can be applied to simplify the boolean output functions of combinational circuits. In this step, the Karnaugh map is applied to simplify and design the output functions of each approximate full adder.
AFA1-3, AFA5 and AFA8 are selected approximate circuits from the previous step. Each of these approximate full adders has two approximate outputs, and for each of the outputs, boolean functions are specified by applying the Karnaugh map. Logic functions of the outputs of AFA1-3, AFA5, and AFA8 are written in (1)-(10), respectively. The output of this step of designing the ICIS full adders is the boolean logic functions of (1)-(10).
@@
STEP 3: The logic functions in (1)-(10) are designed based on boolean logic gates. Applying these logic gates in the design of IMPLY-based circuits is impossible. (1)-(10) should be rewritten so that it is possible to implement them by applying IMPLY and FALSE functions. Implementing all boolean logic functions applying IMPLY and FALSE functions is possible. The implementation of binary logic gates required by (1)-(10) can be investigated and studied in [20]. (1)-(10) are rewritten, applying IMPLY and FALSE functions in (11)-(20).
@@
$\begin{split}[\overline{B_{in}}\to C_{in}]\to[(((B_{in}\to\overline{C_{in}})\to A_{in})\to(\overline{B_{in}\to(C_{in}\to\overline{A_{in}})}))]\end{split}$ (19)
@@
Now, (11)-(20) should be implemented serially by applying a few memristors, three input memristors ($A_{in}$, $B_{in}$, and $C_{in}$), and a maximum of two work memristors ($S_{1}$ and $S_{2}$). The serial implementation algorithm of AFA2, AFA3, and AFA5 are written in Tables 6-8, respectively. These three approximate full adders can be implemented in 6 computational steps according to the presented algorithms and by applying four memristors, including three input memristors and a work memristor.
@@ 6 $A_{in}^{{}^{\prime}}\to B_{in}=B_{in}^{{}^{\prime}}$ $Sum=\overline{C_{out}}$
The implementation algorithm of AFA1 and AFA8 is written serially by applying IMPLY and FALSE functions according to the output functions of these two approximate full adders, (11), (12), (19), and (20). These approximate full adders are removed from the set of ICIS full adders candidates. The main reason for this decision is that according to the complexity of these two circuits and their implementation algorithms, at least five memristors are needed (three input memristors and at least two work memristors) to implement them in 17 computational steps serially. Comparing the computational steps required by these two circuits and their ED with AFA2, AFA3, and AFA5, it can be concluded that the accuracy of computations applying these two circuits will not increase significantly. In contrast, the number of computational steps increases by 65% compared to AFA2, AFA3, and AFA5. So, there would be no justification for presenting these circuits as ICIS approximate full adders.
@@ 3.2 Exact Carry, Inexact Sum IMPLY-based approximate full adder (ECIS)
Table 9: The truth tables of exact full adder and ECIS.
@@
In Table 10, the serial implementation algorithm of ECIS is written in 12 steps by applying five memristors ($A_{in}$, $B_{in}$, $C_{in}$, $S_{1}$ and $S_{2}$). In this algorithm, $Sum$ is calculated in the tenth step, and $C_{out}$ is calculated in the twelfth step.
@@ 3.3 Summary of the proposed approximate full adders
This subsection summarizes the properties of the proposed approximate full adders introduced in subsections 3.1 and 3.2. The main features of ICIS1-3 and ECIS, including error analysis metrics, number of computational steps, and required number of memristors, are written in Table 11.
@@ 4 Simulation results and comparison
In this section, the proposed approximate full adders and SOA [3, 20, 37, 31] are compared with each other by circuit evaluation criteria (subsection 4.1), error analysis metrics (subsection 4.2), and image quality metrics (subsection 4.3). Then, in subsection 4.4, the results of these three simulations are compared with each other by a $FOM$.
@@ 4.1 Circuit-level simulation and analysis
The hardware complexity of the proposed circuits is reduced by applying approximate computing. Approximate full adders’ circuit-level simulation results (energy consumption, computational steps, and estimation of area dissipation) compared to the exact [20, 37] and approximate full adders [3, 31] are critical in determining the improvement of circuit evaluation criteria. Accordingly, it is possible to calculate the improvement of circuit evaluation criteria and apply the simulation results to analyze the application of the proposed circuits in larger computing structures. LTSPICE simulator and Voltage ThrEshold Adaptive Memristor (VTEAM) model are applied for circuit-level simulation of the proposed circuits and SOA [3, 20, 37, 31]. The semi-serial architecture’s compatibility with the crossbar array is moderate [41]; hence, the proposed circuits designed based on serial architecture are not comparable with the semi-serial AFA [32]. In Table 12, the parameters of the applied memristor model and the required parameters for simulating the IMPLY function are written.
@@
$k_{on}$0.5 $\frac{nm}{s}$ $k_{off}$ 1 $\frac{cm}{s}$ $w_{off}$ 0 nm
$w_{on}$ 3 nm $w_{C}$ 107 $pm$ $a_{off}$ 3 $nm$
$a_{on}$ 0 $nm$ $v_{set}$ 1 V $v_{reset}$ 1 V
@@
The proposed circuits are simulated by applying the parameters reported in Table 12 and all the input patterns considered. According to the simulation results, the proposed algorithms for implementing ICIS1-3 and ECIS led to correct results in all cases. Two output waveforms of each of the circuits, including inputs that lead to approximate and exact outputs, as showcased, are shown in Figures 6-9. In Figures 6-9, logic ‘1’ corresponds to LRS, and logic ‘0’ corresponds to HRS.
ICIS1-3 calculate the $Sum$ and $C_{out}$ outputs in 6 computational steps. Each computational step is considered 30 $\mu$s. The $C_{out}$ and $Sum$ outputs of each ICIS1-3 are computed in the fourth (90-120 $\mu$s) and sixth (150-180 $\mu$s) computational steps. Figures 6-8 show the output waveforms of the ICIS1-3. In each of Figures 6-8, two different inputs and corresponding outputs of each input can be seen.
The output waveforms of the ECIS for two input states can be seen in Figure 9. The inputs of $A_{in}B_{in}C_{in}$=“000” and $A_{in}B_{in}C_{in}$=“101” applied to ECIS, its implementation algorithm ran, and the output waveforms depicted in Figure 9. In this circuit, $Sum$ is stored in memristor ‘$C_{in}$’ in the tenth step (270-300 $\mu$s), and $C_{out}$ is stored in memristor ‘$B_{in}$’ in the twelfth step (330-360 $\mu$s).
Figure 6: ICIS1’s waveforms: (a) $A_{in}B_{in}C_{in}$=``$001"$, and (b) $A_{in}B_{in}C_{in}$=``$110"$.
Figure 7: ICIS2’s waveforms: (a) $A_{in}B_{in}C_{in}$=``$000"$, and (b) $A_{in}B_{in}C_{in}$=``$110"$.
Figure 8: ICIS3’s waveforms: (a) $A_{in}B_{in}C_{in}$=``$010"$, and (b) $A_{in}B_{in}C_{in}$=``$111"$.
Figure 9: ECIS’s waveforms: (a) $A_{in}B_{in}C_{in}$=``$000"$, and (b) $A_{in}B_{in}C_{in}$=``$101"$.
Approximate full adders can be applied in n-bit approximate adders in different ways. Generally, approximate full adders are placed in the LSBs, and the exact full adders are applied in the MSBs to control the accuracy of computations. An approximate 8-bit Ripple Carry Adder (RCA) structure is shown in Figure 10. In this structure, 5 LSBs are computed by approximate full adders, and exact full adders calculate 3 MSBs. Three scenarios are considered to report and compare the results of different simulations in this article. The structure in which the mentioned scenarios are examined is the 8-bit approximate RCA. Three, four, and five LSBs of the 8-bit RCA structure are calculated by the approximate full adders in the first, second, and third scenarios, respectively. In the approximate RCA structures of scenarios 1-3, the MSBs (5, 4, and 3, respectively) are calculated by the exact full adders.
@@
The number of computational steps improved by 2%-17% compared to the SIAFA1-4 [3] and SAFAN [31] and 27%-48% compared to 8-bit exact adders by applying ICIS1-3 in scenarios 1-3 (See (25)). The number of computational steps increases 4%-24% when the ECIS is applied in the structures of scenarios 1-3 instead of ICIS1-3, SIAFA1-4 [3], and SAFAN [31]. However, two things should be noted: (1) the accuracy of computations of ECIS is higher than ICIS1-3, SIAFA1-4 [3], and SAFAN [31], and (2) the number of computational steps when ECIS is placed in the 8-bit adders of scenarios 1-3 is reduced 17%-32% compared to the 8-bit exact adders [20, 37].
Energy consumption is one of the most critical circuit analysis metrics. One of the main goals of applying approximate computing in the design of arithmetic circuits is to reduce energy consumption. The LTSPICE energy calculation tool is applied to calculate the energy consumption of the proposed approximate full adders and SOA [3, 20, 37, 31]. First, the sum of the energy consumption of the memristors involved in the computations is measured for each input state of the full adder cell [3]. The average energy consumption of all the input states of the full adder cell is reported as the energy consumption estimation of the full adder cell in Table 14 [3]. According to the results of Table 14, ICIS1-3 have almost equal energy consumption. Based on their implementation algorithms, these three circuits require four memristors to compute their outputs in six computational steps. Their energy consumption is almost equal because of the same number of computational steps and required memristors. The energy consumption of each exact or approximate full adder can be compared with other full adders based on the number of computational steps and the number of memristors required.
@@
(exact and approximate) (×$\times 10^{-9}J$) over [37]
[…]
@@
[…]
According to the results reported in Table 14, the order of energy consumption between the proposed approximate full adders and SOA [3, 31] is as follows: ICIS2 and ICIS3, ICIS1, SAFAN [31], SIAFA4 [3], SIAFA1 [3], SIAFA3 [3], SIAFA2 [3], and ECIS.
The energy consumption improvement of the proposed approximate full adders compared to the exact full adder [37] is written in Table 14. ICIS1-3 improved the energy consumption by 20%-75% compared to the SOA [3, 20, 37, 31]. The energy consumption of the ECIS is higher than the other approximate full adders (ICIS 1-3, SAFAN [31], and SIAFA1-4 [3]). However, this approximate full adder increases computations’ accuracy compared to the others. The ECIS improved the energy consumption of exact full adders [20, 37] by 46% and 49%.

### v2

@@ 3 Proposed IMPLY-based approximate full adders
In this section, the proposed circuits and their design method are described. In subsection 3.1, three approximate full adders are introduced to reduce the hardware complexity, along with an acceptable decrease in computation accuracy. In subsection 3.2, the primary purpose of proposing an approximate full adder is to avoid inexact carry propagation from LSBs to MSBs in an n-bit approximate adder structure.
@@ 3.1 Inexact Carry, Inexact Sum IMPLY-based approximate full adders (ICIS)
Three ICIS full adders (ICIS1-3) are designed in three steps and based on redesigning approximate logic from exact logic by changing the exact full adder’s truth table. It is necessary to determine the conditions of acceptability of approximate full adders’ accuracy before explaining the design steps of the proposed circuits. The maximum value of $ER_{Sum}$ and $ER_{C_{out}}$ should be $\frac{3}{8}$ and $\frac{1}{8}$ to have acceptable output results based on the error analysis metrics of state-of-the-art [3, 4, 15, 17, 18, 21, 32] in error-tolerant applications, e.g., image processing. According to these ERs, the maximum acceptable value of ED for the design of the proposed approximate full adders is also considered ED of 3. Determining these conditions for designing ICIS1-3 is done because the reduction of computation’s accuracy should be limited. In addition to the error evaluation criteria, the computational steps should be significantly reduced compared to the exact full adders. The number of computational steps affects energy consumption directly.
The design steps of ICIS1-3, according to the mentioned design constraints, are as follows:
@@
STEP 1: First, the truth table of the exact full adder is assessed (see Table 4). Then, according to the design constraints mentioned above, the $C_{out}$ of the exact full adder is inverted only in the first state ($A_{in}B_{in}C_{in}$=“000”), and $ER_{C_{out}}$ is $\frac{1}{8}$. In IMPLY logic, it is possible to invert an output in only two computational steps. In the first step, a memristor resets; in the second cycle, the IMPLY function is performed between the output memristor and the one reset in the last cycle. The $Sum$ output of these approximate full adders is assumed to be $\overline{C_{out}}$ to reduce the hardware complexity and the number of computational steps. This process is repeated for states 2 ($A_{in}B_{in}C_{in}$=“001”) to state 8 ($A_{in}B_{in}C_{in}$=“111”) of the exact full adder. Eight approximate full adders’ truth tables are designed by applying this method. We call these eight Approximate Full Adders AFA1-8. These eight approximate full adders’ truth tables can be seen next to the exact full adder’s truth table in Table 4. The truth tables of approximate full adders display exact outputs with ✓ and inexact ones with a ✕. ED of each AFA specified in Table 4. $ER_{C_{out}}$ of these eight approximate full adders is $\frac{1}{8}$. The $ER_{Sum}$ of AFA2-7 is $\frac{3}{8}$, and the $ER_{Sum}$ of AFA1 and AFA8 equals $\frac{1}{8}$. It should be noted that the ED of all these approximate full adders equals 3. Among the AFA1-8 (see Table 4), the truth tables of AFA4, AFA6, and AFA7 are similar to those of SIAFA3, SIAFA1, and SIAFA4 introduced and implemented in our previous research [3]. So, AFA1-3, 5, and 8 are considered implementation candidates according to the error analysis metrics, and the truth table of these approximate full adders is not the same as the truth tables of SIAFA1, SIAFA3, and SIAFA4 [3].
@@
SumSum $C_{out}$ SumSum $C_{out}$ SumSum $C_{out}$ SumSum $C_{out}$ SumSum $C_{out}$ SumSum $C_{out}$ SumSum $C_{out}$ SumSum $C_{out}$ SumSum $C_{out}$
@@
In an alternative technique the truth table of eight approximate full adders is generated by inverting one bit of the exact full adder’s $Sum$ output from state $A_{in}B_{in}C_{in}$=“000” to state $A_{in}B_{in}C_{in}$=“111” similar to what was done for exact $C_{out}$. This time, $C_{out}$ equals $\overline{Sum}$, and $C_{out}$ can be calculated with only two computational steps. The truth tables of Approximate Full Adders 9-16, AFA9-16, are tabulated in Table 5. According to this design method, the $ER_{Sum}$ of these approximate full adders equals $\frac{1}{8}$. As in Table 4, the ED of AFA9-16 is written in the last row of Table 5. AFA9 and AFA16 have acceptable error analysis metrics (See Table 5) based on the conditions specified about the ER and ED of approximate full adders. According to the truth tables of these two circuits, it can be concluded that AFA9 and AFA16 are logically equivalent to AFA1 and AFA8.
@@
SumSum $C_{out}$ SumSum $C_{out}$ SumSum $C_{out}$ SumSum $C_{out}$ SumSum $C_{out}$ SumSum $C_{out}$ SumSum $C_{out}$ SumSum $C_{out}$
@@
STEP 2: After determining the truth table of the acceptable circuits from STEP 1, the Boolean logic function of the outputs of each circuit must be determined. Based on these functions, an algorithm can be provided for implementing these circuits with IMPLY logic by applying the IMPLY and FALSE functions to memristors. The Karnaugh map is one of the methods that can be applied to simplify the Boolean output functions of combinational circuits. In this step, the Karnaugh map is applied to simplify and design the output functions of each approximate full adder.
AFA1-3, AFA5, and AFA8 are selected approximate circuits from the previous step. Each of these approximate full adders has two approximate outputs, and for each of the outputs, Boolean functions are specified by applying the Karnaugh map. Logic functions of the outputs of AFA1-3, AFA5, and AFA8 are written in (5)-(14), respectively. The output of this step of designing the ICIS full adders is the Boolean logic functions of (5)-(14).
@@
STEP 3: The logic functions in (5)-(14) are designed based on Boolean logic gates. Applying these logic gates in the design of IMPLY-based circuits is impossible. (5)-(14) should be rewritten so that it is possible to implement by applying IMPLY and FALSE functions. Implementing all Boolean logic functions applying IMPLY and FALSE functions is possible. The implementation of binary logic gates required by (5)-(14) can be investigated and studied in [23]. (5)-(14) are rewritten, applying IMPLY and FALSE functions in (15)-(24).
@@
$\begin{split}Sum_{AFA8}=[\overline{B_{in}}\to C_{in}]\to[(((B_{in}\to\overline{C_{in}})\to A_{in})\to(\overline{B_{in}\to(C_{in}\to\overline{A_{in}})}))]\end{split}$ (23)
@@
The method of computing the output of $C_{out_{AFA2}}$ is described as an example to better understand how (15)-(24) are written. To generate the output of $C_{out_{AFA2}}$, direct implementation of the functions of NAND and AND logic gates based on IMPLY and FALSE functions and overlapping of the steps related to the FALSE functions in the implementation of these two logic gates ((25) and (26)) has been applied. As can be seen in (27), it is possible to overlap the two FALSE functions and eliminate their consecutive steps. Therefore, using direct implementation of Boolean logic gates using the IMPLY and FALSE functions and overlapping of the steps is the general method used to compute (15)-(24). The method of implementing basic logic gates using the IMPLY and FALSE functions is explained in [23].
$\overline{A_{in}\cdot B_{in}}\equiv(B_{in}\to(A_{in}\to 0)):\alpha$ (25)
$\overline{C_{in}}\cdot(\overline{A_{in}\cdot B_{in}})\equiv\overline{C_{in}}\cdot\alpha\equiv(\alpha\to((C_{in}\to 0)\to 0))\to 0\equiv(\alpha\to C_{in})\to 0\equiv((B_{in}\to(A_{in}\to 0))\to C_{in})\to 0$ (26)
$C_{out_{AFA2}}=\overline{\overline{C_{in}}\cdot(\overline{B_{in}\cdot A_{in}})}\equiv(\overline{C_{in}}\cdot(\overline{B_{in}\cdot A_{in}}))\to 0\equiv(((B_{in}\to(A_{in}\to 0))\to C_{in})\to 0)\to 0$
$\equiv(B_{in}\to(A_{in}\to 0))\to C_{in}$ (27)
Now, (15)-(24) should be implemented serially based on the structure of Figure 2 by applying a few memristors, three input memristors ($A_{in}$, $B_{in}$, and $C_{in}$), and a maximum of two work memristors ($S_{1}$ and $S_{2}$). The serial implementation algorithms of AFA2, AFA3, and AFA5 are written in Tables 6-8, respectively. These three approximate full adders can be implemented in 6 computational steps according to the presented algorithms and by applying four memristors, including three input memristors and a work memristor.
@@ 6 $A_{in}^{{}^{\prime}}\to B_{in}=B_{in}^{{}^{\prime}}$ $Sum=\overline{C_{out}}$
The implementation algorithm of AFA1 and AFA8 is written serially by applying IMPLY and FALSE functions according to the output functions of these two approximate full adders, (15), (16), (23), and (24). These approximate full adders are removed from the set of ICIS full adders candidates. The main reason for this decision is that, according to the complexity of these two circuits and their implementation algorithms, at least five memristors are needed (three input memristors and at least two work memristors) to implement them in 17 computational steps serially. Comparing the computational steps required by these two circuits and their ED with AFA2, AFA3, and AFA5, it can be concluded that the accuracy of computations applying these two circuits will not increase significantly. In contrast, the number of computational steps increases by 65% compared to AFA2, AFA3, and AFA5. So, there would be no justification for presenting these circuits as ICIS approximate full adders.
@@
Figure 2: Schematic representation of ICIS1, ICIS2, and ICIS3.
@@ 3.2 Exact Carry, Inexact Sum IMPLY-based approximate full adder (ECIS)
Table 9: The truth tables of exact full adder and “ECIS” [31].
@@
In Table 10, the serial implementation algorithm of ECIS is written in 12 steps by applying five memristors ($A_{in}$, $B_{in}$, $C_{in}$, $S_{1}$, and $S_{2}$). In this algorithm, $Sum$ is calculated in the tenth step, and $C_{out}$ is calculated in the twelfth step. The circuit-level implementation of ECIS is depicted in Figure 3.
@@ 3.3 Summary of the proposed approximate full adders
This subsection summarizes the properties of the proposed approximate full adders introduced in subsections 3.1 and 3.2. The main features of ICIS1-3 and ECIS, including error analysis metrics, number of computational steps, and required number of memristors, are presented in Table 11.
@@ 4 Simulation results and comparison
In this section, the proposed approximate full adders and state-of-the-art [3, 23, 40, 34] are compared with each other by circuit evaluation criteria (subsection 4.1), error analysis metrics (subsection 4.2), and image quality metrics (subsection 4.3). Then, in subsection 4.4, the results of these three simulations are compared with each other by two $FOMs$.
@@ 4.1 Circuit-level simulation and analysis
The hardware complexity of the proposed circuits is reduced by applying approximate computing. Approximate full adders’ circuit-level simulation results (energy consumption, computational steps, and estimation of area dissipation) compared to the exact [23, 40] and approximate full adders [3, 34] are critical in determining the improvement of circuit evaluation criteria. Accordingly, it is possible to calculate the improvement of circuit evaluation criteria and apply the simulation results to analyze the application of the proposed circuits in larger computing structures. LTSPICE simulator and Voltage ThrEshold Adaptive Memristor (VTEAM) model are applied for circuit-level simulation of the proposed circuits, and state-of-the-art [3, 23, 40, 34]. The semi-serial architecture’s compatibility with the crossbar array is moderate [41]; hence, the proposed circuits designed based on serial architecture are not comparable with the semi-serial AFA [35]. In Table 12, the parameters of the applied memristor model and the required parameters for simulating the IMPLY function are written.
@@
$k_{on}$ -0.5 $\frac{nm}{s}$ $k_{off}$ 1 $\frac{cm}{s}$ $w_{off}$ 0 nm
$w_{on}$ 3 nm $w_{C}$ 107 pmpm $a_{off}$ 3 nmnm
$a_{on}$ 0 nmnm $v_{set}$ 1 V $v_{reset}$ 1 V
@@
The proposed circuits are simulated by applying the parameters reported in Table 12 and all the input patterns considered. According to the simulation results, the proposed algorithms for implementing ICIS1-3 and ECIS led to correct results in all cases. Two output waveforms of each of the circuits, including inputs that lead to approximate and exact outputs, as showcased, are shown in Figures 4-7. In Figures 4-7, logic ‘1’ corresponds to LRS, and logic ‘0’ corresponds to HRS.
[…]
Figure 4: ICIS1’s waveforms: (a) $A_{in}B_{in}C_{in}$=001``001", and (b) $A_{in}B_{in}C_{in}$=110``110".
Figure 5: ICIS2’s waveforms: (a) $A_{in}B_{in}C_{in}$=000``000", and (b) $A_{in}B_{in}C_{in}$=110``110".
Figure 6: ICIS3’s waveforms: (a) $A_{in}B_{in}C_{in}$=010``010", and (b) $A_{in}B_{in}C_{in}$=111``111".
Figure 7: ECIS’s waveforms: (a) $A_{in}B_{in}C_{in}$=000``000", and (b) $A_{in}B_{in}C_{in}$=101``101".
Approximate full adders can be applied to n-bit approximate adders in different ways. Generally, approximate full adders are placed in the LSBs, and the exact full adders are applied in the MSBs to control the accuracy of computations. An approximate 8-bit Ripple Carry Adder (RCA) structure is shown in Figure 8. In this structure, 5 LSBs are computed by approximate full adders, and exact full adders calculate 3 MSBs. Three scenarios are considered to report and compare the results of different simulations in this article. The structure in which the mentioned scenarios are examined is the 8-bit approximate RCA. Three, four, and five LSBs of the 8-bit RCA structure are calculated by the approximate full adders in the first, second, and third scenarios, respectively. In the approximate RCA structures of scenarios 1-3, the MSBs (5, 4, and 3, respectively) are calculated by the exact full adders.
@@
The number of computational steps improved by 2%-17% compared to the SIAFA1-4 [3] and SAFAN [34] and 27%-48% compared to 8-bit exact adders by applying ICIS1-3 in scenarios 1-3 (See (32)). The number of computational steps increases 4%-24% when the ECIS is applied in the structures of scenarios 1-3 instead of ICIS1-3, SIAFA1-4 [3], and SAFAN [34]. However, two things should be noted: (1) the accuracy of computations of ECIS is higher than ICIS1-3, SIAFA1-4 [3], and SAFAN [34], and (2) the number of computational steps when ECIS is placed in the 8-bit adders of scenarios 1-3 is reduced by 17%-32% compared to the 8-bit exact adders [23, 40].
Energy consumption is one of the most critical circuit analysis metrics. One of the main goals of applying approximate computing in the design of arithmetic circuits is to reduce energy consumption. The LTSPICE energy calculation tool is applied to calculate the energy consumption of the proposed approximate full adders, and state-of-the-art [3, 23, 40, 34]. First, the sum of the energy consumption of the memristors involved in the computations is measured for each input state of the full adder cell [3]. The average energy consumption of all the input states of the full adder cell is reported as the energy consumption estimation of the full adder cell in Table 14 [3]. According to the results of Table 14, ICIS1-3 have almost equal energy consumption. The number of computational steps and memristor switching, considering dynamic energy consumption, directly affect the amount of energy consumed in the stateful memristive arithmetic structures. Based on their implementation algorithms, these three circuits require four memristors to compute their outputs in six computational steps. Their energy consumption is almost equal because of the same number of computational steps and required memristors. The energy consumption of each exact or approximate full adder can be compared with other full adders based on the number of computational steps and the number of memristors required.
@@
(exact and approximate) ($\times 10^{-9}J$) over [40]
Exact 1 [23] 1.90859 5 $\%$
@@
SIAFA1 [3] 0.67221 67 $\%$
SIAFA2 [3] 0.86032 57 $\%$
SIAFA3 [3] 0.67221 67 $\%$
SIAFA4 [3] 0.67086 67 $\%$
SAFAN [34] 0.64282 68 $\%$
ICIS1 0.50709 75 $\%$
ICIS2 0.50705 75 $\%$
ICIS3 0.50705 75 $\%$
ECIS 1.02631 49 $\%$
The energy consumption improvement of the proposed approximate full adders compared to the exact full adder [40] is written in Table 14. ICIS1-3, as the most energy-efficient approximate full adders, improved the energy consumption by 20%-75% compared to the state-of-the-art [3, 23, 40, 34]. The energy consumption of the ECIS is higher than of the other approximate full adders (ICIS 1-3, SAFAN [34], and SIAFA1-4 [3]). However, this approximate full adder increases computations’ accuracy compared to the others. The ECIS improved the energy consumption of exact full adders [23, 40] by 46% and 49%.
@@
The limitations of the analyses of this study, along with the functionality assessment of proposed approximate full adders in a 1T1R crossbar structure, are briefly discussed at the end of this subsection. Employing practical models that incorporate various design details, including diverse parasitic features, enhances the quality of the simulation. The authors of this article did not have access to a model based on a specific fabrication technology that accounts for parasitic features in detail, nor to simulation and layout design tools based on these technologies. Accordingly, the model applied in [3, 11, 15, 25, 34, 35, 42, 43] was employed for the simulation to enable a fair comparison between previously verified circuits and the proposed circuits in this article. The authors have also simulated the proposed approximate full adders algorithms in the 1T1R structure in addition to the 1R structure. The main reason for this simulation is to assess the correct functionality of the proposed circuit in this architecture. The use of 1T1R crossbar array architecture can serve as an efficient solution to address design challenges based on 1R crossbar array structures, such as sneak current path [11]. Voltages proportional to the time intervals of the pulse time for the proper functionality of the circuits based on the proposed algorithms have been applied to the bit and word lines. The output memristors in this structure have also been specified in accordance with the proposed algorithm.